OpenStudy (anonymous):

Solve. |x| + 5 = 8 A. {–13} B. {13, –3} C. {–3, 3} D. {–13, –3}

OpenStudy (anonymous):

lxl +5=8 the integer solution is x = ±3

OpenStudy (anonymous):

soooo...... can you guess what the answer is? :)

OpenStudy (anonymous):

x=+3

OpenStudy (anonymous):

for that equation you go x= 8-5 x=3 x= +3 because its only trying to find out what one number is, x+5=8 is the same as saying x+_=8 so you work it out the same way!